OpenStudy (anonymous):

what is sqrt(7+4sqrt3)+sqrt(7-4sqrt3)?thnx!

OpenStudy (anonymous):

do you want to find the numerical value of this?

OpenStudy (anonymous):

yes pls. and the solution on how to get it?

OpenStudy (anonymous):

u can of couse do this with a calculator

OpenStudy (anonymous):

but our teacher did not allow us to use calculatro

OpenStudy (anonymous):

i cant see a way to simplify it i'm afraid of course i could be wrong

OpenStudy (anonymous):

oh ok. thnx anyway!:-)))

OpenStudy (anonymous):

sorry i couldn't help

OpenStudy (anonymous):

well its ok! thnx!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

You can use this identity \[(a+b)(a-b) = a^2 - b^2\] So we get \[(7+4\sqrt{3})(7-4\sqrt{3}) = 7^2- (4\sqrt{3})^2\] \[49- 16(3) =49-48=1\]

OpenStudy (anonymous):

@rnpua , hope this is what you were looking for :)

OpenStudy (anonymous):

right - i missed that one!!

OpenStudy (anonymous):

yeah, i tried that, but i realized it was addition, not multiplication though.haha.but thnx @shivam_bhalla!